### v3

# Towards General Purpose Vision Foundation Models for Medical Image Analysis: An Experimental Study of DINOv2 on Radiology Benchmarks

Abstract
The integration of deep learning systems into healthcare has been hindered by the resource-intensive process of data annotation and the inability of these systems to generalize to different data distributions. Foundation models, which are models pre-trained on large datasets, have emerged as a solution to reduce reliance on annotated data and enhance model generalizability and robustness. DINOv2 is an open-source foundation model pre-trained with self-supervised learning on 142 million curated natural images that exhibits promising capabilities across various vision tasks. Nevertheless, a critical question remains unanswered regarding DINOv2’s adaptability to radiological imaging, and whether its features are sufficiently general to benefit radiology image analysis. Therefore, this study comprehensively evaluates DINOv2 for radiology, conducting over 100 experiments across diverse modalities (X-ray, CT, and MRI). To measure the effectiveness and generalizability of DINOv2’s feature representations, we analyze the model across medical image analysis tasks including disease classification and organ segmentation on both 2D and 3D images, and under different settings like kNN, few-shot learning, linear-probing, end-to-end fine-tuning, and parameter-efficient fine-tuning. Comparative analyses with established supervised, self-supervised, and weakly-supervised models reveal DINOv2’s superior performance and cross-task generalizability. The findings contribute insights to potential avenues for optimizing pre-training strategies for medical imaging and enhancing the broader understanding of DINOv2’s role in bridging the gap between natural and radiological image analysis. Our code is available at https://github.com/MohammedSB/DINOv2ForRadiology
keywords:
Representation Learning, Self-Supervised Learning, Foundation Models, Computer Vision, Medical Image Analysis.
journal: Medical Image Analysis
654321654321DINOv2ViT-gDINOv2ViT-LOpenCLIPCLIPMAEMSNSegmentation RankClassification RankRelative Performance for Classification and Segmentation
Figure 1: Cross-task generalizability of DINOv2 compared to other models. The horizontal axis shows the relative performance or ranking of each of the models on segmentation tasks, while the vertical axis does the same for classification tasks. The ranking is calculated as the average ranking across all segmentation or classification tasks, where rank 1 means the model performs the best relative to the other models. Models pre-trained with weakly-supervised learning perform well only on classification tasks, while MAE performs well only on segmentation. DINOv2 can generalize across both tasks and outperforms all other models for classification.
1 Introduction
Radiology imaging plays a pivotal role in modern medicine, serving as an indispensable tool for accurate and timely diagnosis [1]. The importance of radiological imaging lies in its ability to provide detailed and non-invasive visualizations [2] of the internal structures and functions of the human body. Deep learning-based computer vision methods have been successful at analyzing and processing radiological imaging, leading to systems that can extract clinically-relevant information with high accuracy [3]. However, the success of these systems has been reliant on annotated medical data, which is expensive to obtain because it requires the time and effort of trained radiologists [4]. Moreover, these systems can achieve high-level accuracy only within a specified scope, but fail to generalize across domains, tasks, and slight data distribution shifts [5].
Recently, the field of computer vision has seen a rise in interest for general-purpose models that are optimized to function across different tasks and domains [6, 7, 8, 9]. These models, grouped under the term “Foundation Models” (FMs), usually contain parameters ranging from hundreds of millions to tens of billions and are trained on large datasets, on the order of tens of millions. As a result of this large-scale training, these FMs often achieve state-of-the-art (SoTA) results and impressive zero-shot and few-shot performance and generalizability [8, 9]. For these reasons, foundation models have gained traction in deep learning-based medical image analysis research [10, 11, 12, 13, 14], as they hold promise for reducing the reliance on the expensive process of annotating medical data and towards the goal of building generalist medical artificial intelligence systems that can function across a variety of tasks and domains [15].
What Are Foundation Models?
The term foundation model is an umbrella term that covers a wide range of seemingly different models. In the most general sense, foundation models are large models trained on large datasets and can generalize across tasks and/or domains [16]. To make the term more useful in our analysis, however, we group FMs using two methods of categorization. First, we divide FMs depending on their training paradigm into three groups: self-supervised, weakly-supervised, and supervised foundation models. Weakly-supervised and supervised foundation models require correspondence in the training data. In these paradigms, the training data is required to be available in pairs: an X-ray examination and a corresponding interpretation, diagnosis, or segmentation mask for example. Models like OpenAI’s CLIP [7] and Meta’s Segment Anything Model (SAM) [9] fall under this category. Self-supervised foundation models, on the other hand, require one input data to train (an image, text, audio, etc.) Meta’s DINOv2 [8] and Google’s Universal Speech Model (USM) [17] belong to this category.
Additionally, we categorize FMs into two groups depending on the generalizability of their produced representations: general purpose (also called task-agnostic), and task-specific FMs. General purpose foundation models produce features that generalize across more than one task, segmentation and classification for example, while task-specific models specialize on only one task. DINOv2 [8] and USM [17] fall under the former category while SAM [9] is under the latter.
Because of the less restrictive training dependencies for self-supervised FMs, we believe that they are a promising option to explore for future research especially for the medical imaging domain, since public medical image data are abundant but often without corresponding descriptions or diagnoses, either due to patient privacy barriers, data collection overhead, or other issues. Moreover, FMs that produce representations that can be used across a variety of tasks are desired because it is usually a signal of model robustness. As a result, we focus our attention on self-supervised general-purpose FMs for medical image analysis. Specifically, we adopt Meta’s DINOv2 [8] model, a publicly available general-purpose vision foundation model that can extract robust representations across different vision tasks, for experimentation on a wide range of medical disease classification and organ segmentation benchmarks, across different radiological exams and under different evaluation settings.
What is DINOv2?
DINOv2 is a successor of DINO [18] and constitutes both a self-supervised pre-training method based on DINO and iBOT [19], and a collection of models pre-trained using that method. It was released to the public by Meta in April 2023 and promises robust representations that enable general-purpose functionality with visual-only data [8]. The released models were pre-trained on a dataset of 142 million carefully curated natural images, called LVD-142M. Roughly 100 million of these are images that are similar to ImageNet, curated from calculating similarly of web-scarpered images with ImageNet21k dataset [20]. The remaining images were retrieved based on their similarity to Caltech 101 [21], ADE20k [22], and Google Landmarks v2 [23], among others. The models achieve competitive performance on classification, segmentation, depth estimation, and image retrieval tasks across both image and video benchmarks. Moreover, because of the adopted discriminative self-distillation approach, DINOv2 performs well “out-of-the-box,” without the need to fine-tune the encoder. The ViT-g/14 version of the model achieves an accuracy of 86.5% and 83.5% on Linear-probing and kNN evaluations, respectively, on ImageNet-1k, outperforming other weakly-supervised and self-supervised methods. This capability to perform well out-of-the-box is appealing, especially in the medical domain, as it implies competitive performance even in low-data and low-computation settings.
Figure 2: PCA component visualization. Following [8], the PCA is computed between patches of images that are in the same column, and the first 3 components are shown. This is done for X-ray, CT, and MRI scans. Thresholding is used on the first component to remove the background.
Contribution
In this paper, we set to work towards vision-only general-purpose foundation models for the medical domain by adopting DINOv2 for disease classification and organ segmentation benchmarks. We perform comprehensive evaluations of DINOv2 across various scenarios for multiple radiology modalities, exploring both low (few-shot) and high data settings spanning X-ray, CT, and MRI examinations.
We evaluate DINOv2 on both disease classification and organ segmentation benchmarks. For disease classification tasks, we evaluate the model on kNN, linear-probing, few-shot learning, parameter-efficient fine-tuning, and end-to-end fine-tuning scenarios. For organ segmentation tasks we compare lightweight and non-lightweight decoders, while we keep the DINOv2 backbone frozen. We compared the results to supervised, weakly-supervised, and self-supervised models. As far as we know, there is no comprehensive analysis of DINOv2 on medical benchmarks that evaluates the model across disease classification and organ segmentation tasks and on different radiological examinations. Our contributions can be summarized as follows:
  • 1.
    Evaluate DINOv2 on disease classification and organ segmentation radiology benchmarks, under different evaluation settings. We conclude that DINOv2 outperforms self-supervised and weakly-supervised models on classification tasks and is competitive with supervised models pre-trained on ImageNet21K. For segmentation tasks, we observe that DINOv2 outperforms weakly-supervised methods by a large margin, and outperforms or is competitive with supervised methods trained end-to-end.
  • 2.
    Analyze the cross-task generalizability of DINOv2 compared to supervised, self-supervised, and weakly-supervised models. Our analysis concludes that weakly-supervised models perform well on classification tasks only, and masked imaged modeling self-supervision performs well only on segmentation tasks, while DINOv2 can generalize across both tasks. DINOv2 also outperforms supervised classification models when evaluated on segmentation tasks, and outperforms SAM for disease classification.
  • 3.
    Employ parameter-efficient fine-tuning with a DINOv2 ViT-g/14 model on large X-ray classification datasets and provide a performance and efficiency comparison of PEFT with end-to-end fine-tuning and linear-probing. We conclude that using PEFT can yield performance that is competitive with end-to-end fine-tuning using less than 1% of the total model parameters.
2 Related Works
Supervised foundation models. Training a large-scale network with supervised learning is non-trivial because it is difficult to combine different labels together. This is even more true in the medical domain where labeled data is scarce. Even still when labels are seemingly similar, other issues can arise. Cohen et al. [24] have attempted to train a large-scale X-ray foundation model with supervised learning by combining public chest X-ray disease classification datasets that share diseases. their analysis highlights the fact that discrepancies in labeling caused by disease concept shift across hospitals and observers can lead to worse generalizations for deep learning models.
Regardless, there successful attempts at building a large-scale supervised foundation model. The Segemnt Anything Model (SAM) [9], a supervised foundation model for prompt-based segmentation trained with automatically extracted segmentation masks, is one of those attempts. Since the release of SAM, its applicability for the medical images has been explored [25, 26, 27]. It’s performance for medical segmentation was found to be inadequate in most cases and highly variable. Because of that, Ma et al. [14] have developed MedSAM by adopting SAM for the medical domain using a dataset of more than one million medical images across diverse segmentation tasks and examination modalities. MedSAM significantly outperforms SAM on medical tasks and achieves comparable performance to U-Net [28] models. Still, MedSAM is limited to medical segmentation, and it is unclear whether it can be generalized to other medical image analysis tasks.
Weakly-supervised foundation models. Previous research has explored weakly-supervised vision-language models for zero-shot classification, visual question answering, and image generation in the medical domain, achieving SoTA or competitive performance on radiology benchmarks [29, 30, 31, 32]. Zhang et al. [31] applied contrastive learning to paired X-ray image-report for learning visual representations. However, the availability of corresponding image-text pairs is a stringent requirement to be applicable for most medical datasets. Moreover, as highlighted by [29], a false negative issue exists when applying this learning paradigm in the medical domain because, unlike diverse natural images, medical reports can describe other patients that are not necessarily paired. Instead, Wang et al. [29] proposed MedCLIP, employing an extension of CLIP’s contrastive pre-training method that utilizes unpaired examination-report samples, using the MIMIC-CXR [33] and CheXpert [34] datasets. They outperform previous SoTA on zero-shot and supervised classification on four radiology benchmarks. Yet, the performance of MedCLIP, along with other vision-language models, is still limited by the availability of text-image data.
Self-supervised foundation models. Self-supervised pre-training applied to medical datasets has recently achieved SoTA on CT and MRI segmentation benchmarks [35, 36]. Tang et al. [36] pre-trained a Swin UNETR [37] architecture with self-supervision on 5,050 CT volumes and outperformed previous SoTA on the BTCV [38] and MSD [39] competitions. However, to the best of our knowledge, there is still no self-supervised general-purpose FM for the medical domain that has achieved consistently competitive results across tasks and radiological examinations.
3 Methodology
In this section, we describe the motivation for adopting DINOv2 for this study and outline the settings under which we performed our experiments. We do not detail the pre-processing and hyperparameter tuning pipeline and instead publicly release our code (available in Section 6), in which these details can be found.
Motivation for Using DINOv2
There are many vision foundation models trained with supervised, self-supervised, and weakly-supervised learning on natural images [7, 9, 6, 40, 41]. However, we decided to employ DINOv2 in our analysis because of the robustness of its representations, achieving competitive performance in multiple downstream tasks, across vision modalities (image and video), and, most importantly, in out-of-the-box evaluations. The DINOv2 training paradigm was specifically designed to generate powerful representations on out-of-the-box kNN evaluations and outperforms many other weakly-supervised and self-supervised foundation models in kNN and linear-probing [8].
Datasets
We evaluated DINOv2 on 9 public radiology benchmarks, spanning X-ray, CT, and MRI examinations (Exam.) for disease classification (CLS, 4 datasets) and organ segmentation (SEG, 5 datasets) tasks. A summary of the used datasets is shown in Table 1. “# Classes” describes the number of classes and “# Images/Volumes” describes the number of images for 2D datasets and volumes for 3D datasets, and not the number in the original dataset. For some of the datasets shown, we only selected a sample of the entire available data. Moreover, some of the datasets used do not have a predefined test set. On these datasets, we used systemic sampling to divide the dataset into train, evaluation, and test subsets. All the datasets used are publicly available, and we provide a link for downloading our specific train, validation, and test split in Section 6. Moreover, The data pre-processing pipeline is also available in the GitHub repository in Section 6.
The datasets were chosen based on the diversity of tasks and modalities and not on their clinical usefulness or value. Classifying different brain tumors types might not be a clinically applicable task, but we still included it in our analysis to gauge how each model performs in disease classification with MRI examinations.
Table 1: The datasets used in our analysis. For datasets that do not have a standardized test set, we chose the test set using systematic sampling. All the datasets and splits are public.
Dataset Exam. Task Labels # Classes # Images/Volumes Dim NIH Chest X-ray [42] X-ray CLS Thorax Diseases 14 112,120 2D CheXpert [34] X-ray CLS Thorax Diseases 5 161,792 2D Montgomery County (MC) [43] X-ray SEG Lung 3 138 2D SARS-CoV-2 [44] CT CLS COVID-19 Diagnosis 2 210 3D AMOS [45] CT SEG Abdominal Organs 15 150 3D MSD Spleen [39] CT SEG Spleen 2 40 3D MSD Hipp [39] MRI SEG Hippocampus Head and Body 3 260 3D MSD Heart [39] MRI SEG Left Atrium 2 20 3D Brain Tumor [46] MRI CLS Tumor Types 3 3,064 2D
Evaluation Settings
In our analysis, we focused mainly on the “out-of-the-box” performance of DINOv2, where we trained classification or segmentation heads while keeping the backbone frozen. This resulted in preferable lightweight training that requires fewer labeled instances, computational resources, and training time. Additionally, we also performed end-to-end fine-tuning and parameter-efficient fine-tuning evaluations for performance comparison to this lightweight training paradigm.
We experimented with the original DINOv2 ViT-g/14 and the three smaller distilled versions (ViT-L/14, ViT-B/14, and ViT-S/14). We compared the performance and generalizability of DINOv2 to supervised, self-supervised, and weakly-supervised models. Table 2 detailed all the models used. Section 3.3.1 further details the evaluation settings for the classification experiments, while 3.3.2 does the same for segmentation settings.
Table 2: Models used in our analysis. Description of the backbones used along with their parameter count and pre-training settings. ”Used For” describes what we used each model for and ”# Images” describes the number of images in the pre-training dataset.
Method Architecture Dataset Used For # Images # Params. Citation Supervised CLS DenseNet201 ImageNet1k CLS 1.3M 20M [47] ResNet152 ImageNet1k CLS 1.3M 60M [48] VGG19 ImageNet1k CLS 1.3M 144M [49] ViT-L/16 ImageNet21k CLS, SEG 14M 300M [50] SAM ViT-L/16 SA-1B CLS, SEG 11M 300M [9] Weakly-Supervised CLIP ViT-L/14 WIT-400M CLS, SEG 400M 300M [7] OpenCLIP ViT-H/14 LAION-2B CLS, SEG 2, [^0] M 632M [41] Self-Supervised MAE ViT-L/16 ImageNet1k CLS, SEG 1.3M 300M [51] MSN ViT-L/16 ImageNet1k CLS, SEG 1.3M 300M [52] DINOv2 ViT-S/14 LVD-142M CLS 142M 21M [8] ViT-B/14 LVD-142M CLS 142M 86M [8] ViT-L/14 LVD-142M CLS, SEG 142M 300M [8] ViT-g/14 LVD-142M CLS, SEG 142M 1,100M [8]
A critical part of our work is that we standardized the training and hyperparameter tuning pipeline for each model and evaluation task, with the goal of isolating model performance. In this way, we can gain a clearer understanding of which models perform best under the same evaluation settings.
Disease Classification
We performed four main types of experiments: kNN, linear-probing, few-shot learning, and fine-tuning. (1) kNN was performed on the normalized features of the last backbone layer. (2) For linear-probing, a single linear layer was attached on top of the backbone. (3) In few-shot learning, we trained a linear layer on top of frozen features. (4) For fine-tuning, we used both parameter-efficient fine-tuning methods like LoRA [53] and BitFit [54] and end-to-end fine-tuning.
When performing linear-probing with ViT architectures, the linear layer takes either the CLS token or the CLS token concatenated with the average of all patch tokens, depending on which method yielded higher performance in the validation set. In multi-labeled classification, we predicted each class as a binary and averaged the model’s result across all classes, and we used the method described by [55] for kNN. For 3D volumes, the embeddings for all slices were averaged before being passed into the classification head.
Organ Segmentation
For organ segmentation evaluations, we kept the encoder frozen and attached a U-Net, hierarchical decoder on top. The U-Net decoder is made up of four blocks, where each block consists of one convolutional layer along with ReLU activation function and batch normalization. Skip connections were obtained from the previous four blocks of the transformer model and concatenated to the features at each U-Net layer, as is done in the classical U-Net architecture [28].
We also compared the performance of DINOv2 to segmentation architectures that are commonly used in medical image analysis including U-Net and TransUnet, trained end-to-end in a supervised fashion, and experimented with using a lightweight single linear layer decoder as a comparison with the U-Net decoder. For processing 3D volumes, we segmented each slice independently.
4 Results
In this section, we report our results across the different evaluation settings, tasks, and radiological modalities. We compare the performance of DINOv2 to weakly-supervised models including CLIP [7] and OpenCLIP[41], and self-supervised models including MAE [51] and MSN [52], and supervised models. We used the area under the operating receiver curve (AUROC) as a performance metric for classification tasks, and the average of the dice and jaccard scores as a metric for segmentation. We report only the best checkpoint across the epochs tested for each model.
In Section 4.1 and we will evaluate DINOv2 on disease classification and organ segmentation tasks. For segmentation, we will also show a comparison between using a linear layer decoder and a U-Net, hierarchical decoder on top of the frozen DINOv2 ViT-L/14 features. Then, in Section 4.3, we will explore the few-shot learning capability of DINOv2 on both disease classification and organ segmentation tasks. After that, in Section 4.4, we perform parameter-efficient fine-tuning to compare the performance and efficiency of using PEFT with end-to-end fine-tuning and linear probing. In Section 4.2, we will analyze the cross-task generalizability of DINOv2 compared to the other models on disease classification and organ segmentation tasks. Finally, 4.5 shows qualitative results of DINOv2 features on X-ray, CT, MRI modalities, and organ segmentation results of linear and U-Net decoders trained on top of frozen DINOv2 ViT-L/14 features.
Disease Classification and Organ Segmentation
We will start by analyzing the performance of DINOv2 ViT-g/14 and ViT-L/14 on linear-probing. Figure 3 shows the AUROC scores of both DINOv2 models compared to weakly-supervised and self-supervised models on four disease classification benchmarks. DINOv2 outperforms all other models on the more difficult large X-ray datasets and achieves performance closer to weakly-supervised models that are trained on much larger datasets. The ViT-L version of the model seemingly underperforms on 3D CT classification but is competitive with the ViT-g/14 version on all other classification tasks. Additionally Table 3 shows the linear-probing performance of all DINOv2 models compared to supervised learning methods on X-ray, CT, and MRI disease classification datasets. DINOv2 performs on par or slightly better compared to supervised methods, and outperforms by a relatively larger margin commonly-used CNN supervised models.
Figure 3: Linear probing for disease classification compared to self-supervised and weakly-supervised methods. The figure shows the performance of linear probing DINOv2 compared to other self-supervised and weakly-supervised models.
Table 3: Linear probing for disease classification compared to supervised methods. The Table compares supervised models with DINOv2 pre-trained models in Linear-probing settings on X-ray, CT, and MRI datasets. All models were tested under the same settings.
Method Architecture NIH Chest X-ray CheXpert SARS-CoV-2 Brain Tumor Supervised DenseNet201 0.735 0.795 0.973 0.960 ResNet152 0.718 0.779 0.936 0.948 VGG19 0.696 0.750 0.891 0.933 ViT-L/16 0.751 0.829 0.983 0.975 DINOv2 ViT-S/14 0.747 0.805 0.943 0.962 ViT-B/14 0.755 0.812 0.922 0.972 ViT-L/14 0.763 0.821 0.950 0.974 ViT-g/14 0.759 0.818 0.978 0.976
Table 4 shows the kNN, linear-probing, and end-to-end fine-tuning results of DINOv2 compared to other supervised, weakly-supervised, and self-supervised methods on the NIH Chest X-ray and CheXpert datasets. Just like in linear probing, DINOv2 outperforms self-supervised methods in end-to-end fine-tuning but is competitive with weakly-supervised and supervised methods. Also important to highlight that DINOv2 under-performs on kNN evaluations compared to other methods, even though its features were designed to maximize kNN results. This might be explained by the domain shift between natural images in the pre-training and medical images, making the out-of-the-box kNN evaluations more random.
Table 4: DINOv2 performance comparison on large X-ray datasets. DINOv2 outperforms other methods on Linear-probing and fine-tuning but under performs on kNN evaluations. Multiple learning rates for tuning each backbones were tested, and only the best is shown.
NIH Chest X-ray CheXpert Method Architecture kNN Linear-probing Fine-tuning kNN Linear-probing Fine-tuning Supervised DenseNet201 0.675 0.735 0.769 0.783 0.795 0.882 Resnet152 0.668 0.718 0.752 0.766 0.779 0.868 VGG19 0.644 0.696 0.711 0.728 0.750 0.870 ViT-L/16 0.663 0.751 0.761 0.777 0.829 0.873 MAE ViT-L/16 0.659 0.724 0.743 0.785 0.789 0.821 MSN ViT-L/16 0.692 0.742 0.707 0.807 0.753 0.802 CLIP ViT-L/14 0.655 0.739 0.697 0.742 0.816 0.842 OpenCLIP ViT-H/14 0.659 0.737 0.770 0.744 0.814 0.847 DINOv2 ViT-L/14 0.663 0.763 0.717 0.771 0.821 0.786 ViT-g/14 0.659 0.759 0.769 0.768 0.818 0.848
Table 5: DINOv2 on organ segmentation. A comparison between using a frozen DINOv2 backbone and other commonly-used segmentation models initialized from scratch.
Method Architecture MC AMOS MSD Heart MSD Hipp MSD Spleen MAE ViT-L/16 0.969 0.547 0.864 0.799 0.853 MSN ViT-L/16 0.961 0.407 0.810 0.788 0.784 CLIP ViT-L/14 0.955 0.445 0.762 0.748 0.747 OpenCLIP ViT-H/14 0.962 0.501 0.793 0.779 0.803 DINOv2 ViT-L/14224 0.966 0.512 0.792 0.812 0.813 ViT-L/14448 0.974 0.592 0.869 0.789 0.898 ViT-g/14224 0.966 0.511 0.804 0.761 0.802 ViT-g/14448 0.973 0.642 0.875 0.729 0.900
Moreover, we provide a comparison for segmentation tasks in Table 5. When comparing just image sizes of 224x224, DINOv2 performs the second-best, just behind MAE, which outperforms DINOv2 on all but one segmentation task. This can be explained given MAE’s inherently spatial learning objective. An important point to note here is that all models except DINOv2 were pre-trained on image sizes of 224x224. At the end of the DINOv2 pre-training, it was adapted to image size 518x518, leading to worse performance on smaller image sizes due to positional encoding interpolation [8]. Even then, DINOv2 still outperforms weakly-supervised models and MSN pre-training. When using image sizes closer to DINOv2’s pre-training, it outperforms all other methods. However, it is difficult to determine how much of this increased performance is due to reduced interpolation of positional encoding or to higher details resulting from the larger images.
Table 6: DINOv2 compared to segmentation architectures. A comparison between using a frozen DINOv2 backbone and other commonly-used segmentation models initialized from scratch. The parameter count given is when there is one output class.
Method Architecture Trainable Params. Total Params. MC AMOS MSD Heart MSD Hipp MSD Spleen Scratch U-Net 31M (100%) 31M 0.973 0.432 0.911 0.593 0.826 TransUnet 324M (100%) 324M 0.974 0.535 0.892 0.821 0.855 DINOv2 ViT-L/14 17M (5%) 317M 0.974 0.592 0.869 0.789 0.898 ViT-g/14 38M (3%) 1,138M 0.973 0.642 0.875 0.729 0.900
DINOv2 compared to segmentation architectures. In Table 6 we compare U-Net decoders on top of frozen DINOv2 ViT-L/14 and ViT-g/14 features with U-Net and TransUnet models trained end-to-end from scratch. The results of all models are similar on the easier MC dataset, but DINOv2 outperforms the other U-Net and TransUnet on the more difficult AMOS multi-organ segmentation task even with a frozen encoder and less trainable parameters. On the MSD datasets, DINOv2 is competitive but there is variability in the performance.
U-Net vs linear decoders. Table 4 shows a performance comparison between using a linear layer decoder and a U-Net decoder on top of frozen DINOv2 ViT-L/14 features on the organ segmentation datasets. The linear layer decoder evaluations are used to isolate the performance of the DINOv2 encoder, analogous to the purpose of kNN classification evaluations. On the MC dataset where the target mask is large, linear and U-Net performance is comparable, highlighting the strong out-of-the-box features of DINOv2. The gap in performance increases, however, on MSD datasets where target masks are usually much smaller, making them harder to predict with a single layer. Figure 6 shows qualitative segmentation for both methods.
Cross-task Generalizability Analysis
In this section, we evaluate whether DINOv2 can produce representations that are more generalizable across tasks, compared to other models. To accomplish this task, we plot in Figure 1 the relative performance of weakly-supervised and self-supervised methods across both disease classification and organ segmentation. The horizontal axis shows the relative performance or ranking of each of the models on segmentation tasks, while the vertical axis does the same for classification tasks. The ranking is calculated as the average ranking across all segmentation or classification tasks, where rank 1 means the model performs the best relative to the other models. An interesting observation is that models pre-trained with weakly-supervised learning perform well only on classification tasks, while MAE performs well only on segmentation. DINOv2 can generalize across both tasks and outperforms all other models for classification.
DINOv2 corss-task generalizability compared to supervised methods. Additionally, we carry out two experiments to compare the task-generalizability of DINOv2 with supervised methods. First, we compare the segmentation performance of a DINOv2 pre-trained ViT-L/14 with a ViT-L/16 pre-trained with supervised learning on ImageNet21k. Table 7 shows the results. DINOv2 outperforms the supervised ViT on 4 out of the 5 organ segmentation tasks, especially in the challenging AMOS multi-organ segmentation task.
Our second experiment compares the performance of the SAM image encoder with DINOv2 on classification tasks. SAM was trained for prompt-based segmentation and does not have a CLS token. To perform classification with SAM we averaged all the patch embeddings and treated the result as a CLS token. Table 8 shows the results. Images are of size 1024x1024 and only a subset of each dataset was used because of computational limits. DINOv2 significantly outperforms SAM on both datasets, highlighting the cross-task generalizability of the model.
Table 7: DINOv2 vs. ImageNet21k pre-trained ViT on organ segmentation. DINOv2 outperforms the supervised pre-trained ViT on 4 of the 5 tasks.
Method Architecture Image Size Montgomery County AMOS MSD Heart MSD Hipp MSD Spleen Supervised ViT-L/16 224 0.963 0.433 0.825 0.750 0.773 DINOv2 ViT-L/14 224 0.966 0.512 0.792 0.812 0.813 ViT-L/14 448 0.974 0.592 0.869 0.789 0.898
Table 8: SAM vs DINOv2 on X-ray on disease classification. Only a subset of the entire dataset was and all the images are of size 1024x1024. The average of patch embeddings were used as a CLS token for SAM.
Method Architecture NIH Chest X-ray CheXpert SAM ViT-L/16 0.714 0.792 DINOv2 ViT-L/14 0.755 0.816
MCAMOSMSD HeartMSD HippMSD Spleen0.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111+2.9+7.7+16.6+12.5+18.2DatasetAveraged Dice and JaccardLinearU-Net
Figure 4: Linear vs. U-Net decoder. Comparison of the Linear and U-Net decoders on the four segmentation tasks used in this analysis. The backbones used in both is a DINOv2 ViT-L/14.
Few-shot Learning
To measure DINOv2’s ability to adapt to new distributions using a few labeled instances, we perform few-shot learning for both disease classification and organ segmentation on X-ray datasets.
At the top row of Figure 5, we start by comparing DINOv2 ViT-L/14 to weakly-supervised and self-supervised methods. The top-left subplot shows the performance on the NIH Chest X-ray disease classification dataset, while the top-right subplot shows the performance on the MC lung segmentation datasets. For disease classification, there is no clear trend when the number of patients used for each class is between 1 and 4, but when 8 patients are used, DINOv2 outperforms all other methods. For organ segmentation, DINOv2 outperforms all other methods from the start and is only worse than MAE when the entire dataset is used.
A similar trend can be observed at the bottom row of Figure 5, where we compare DINOv2 with supervised methods. The bottom-left subplot shows that DINOv2 outperforms other methods when the number of patients are 8 or more, but to a lesser degree compared to self-supervised and weakly-supervised methods. For organ segmentation, DINOv2 outperforms other models by a large margin when using less than eight instances, which is somewhat expected given it was pre-trained while the other segmentation models were not.
124816All0.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.8NIH Chest X-ray datasetDINOv2OpenCLIPCLIPMAEMSN
124816All0.90.90.90.91111Montgomery County datasetDINOv2OpenCLIPCLIPMAEMSN
124816All0.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.8N patientsDINOv2ViT21kDenNet1kResNet1kVGG1k
124816All0.40.40.40.40.60.60.60.60.80.80.80.81111N ImagesDINOv2U-NetTransUnet
Figure 5: Few-shot disease classification and organ segmentation. The top row compares DINOv2 ViT-L/14 with weakly-supervised and self-supervised methods on the NIH Chest X-ray and MC datasets. The bottom row provides a comparison with supervised methods. For disease classification, there is no clear trend when the number of patients used for each class is between 1 and 4, but when 8 patients are used, DINOv2 clearly outperforms all other methods. For organ segmentation DINOv2 outperforms all other methods from the start.
Parameter-efficient Fine-tuning
We experiment with parameter-efficient fine-tuning (PEFT) techniques on DINOv2 ViT-g/14, which, as a whole, contains 1.1 billion parameters. PEFT methods are used to enable efficient adaptation of large models to downstream tasks, usually achieving performance that is on par with end-to-end fine-tuning while requiring a lot less compute and memory. Previous work by Dutt et al. [56] has highlighted the opportunity of employing PEFT to tune large foundation models for medical image analysis.
We employ two different PEFT techniques: LoRA [53] and BitFit [54]. LoRA is an additive method that inserts trainable decomposition matrices in the layers of a transformer, while BitFit is a selective method that unfreezes only the bias terms of the model. Table 9 shows a result and efficiency comparison between the two PEFT methods with a comparison to end-to-end fine-tuning and linear-probing on the NIH Chest X-ray and CheXpert datasets using the DINOv2 ViT-g/14 model.
Table 9: PEFT on DINOv2 ViT-g/14. Both LoRA and BitFit achieve results that are better than linear-probing while adapting less than 1% of the total parameters.
Method Trainable Params. (%) NIH Chest X-ray CheXpert Fine-tuning 1,100M (100%) 0.769 0.848 Linear-probing 1,500 (1e-6%) 0.759 0.818 LoRA 8M (0.7%) 0.767 0.823 BitFit 0.8M (0.07%) 0.768 0.817
Qualitative Results
In this section we will show qualitative results of DINOv2 features using principal component analysis (PCA) performed on DINOv2 patch features on X-ray, CT, and MRI scans, following the method delineated in [8]. We will also provide organ segmentation results of linear compared U-Net decoders.
PCA visualization. Figure 2 shows the first three PCA components. The PCA is computed between patches of images that are in the same column, and the first 3 components are shown for X-ray, CT, and MRI scans. Thresholding is used on the first PCA component to remove the background. Just like in natural images [8], the colors of the three PCA components correspond well with the same parts of images in the same category. This is an easier task however, compared to natural images, because there is less variability between examinations on medical images compared to natural images.
U-Net and linear decoder visualization. We also show a visualization of linear and U-Net decoders trained on top of DINOv2 ViT-L/14 features. The linear layer decoder performs surprisingly well, but is limited, especially on smaller masks, due to the smaller decoding map (32x32 pixels interpolated to 448x448) and less adjustable parameters. As expected, the U-Net segmentation results are smoother and represents the ground truth mask more accurately, but is still limited due to the frozen encoder.
Image
Ground Truth
Linear
Image
Ground Truth
Linear
U-Net
Figure 6: Linear vs. U-Net visualization. The figure shows a qualitative comparison between the linear layer and the U-Net decoder.
5 Discussion
Foundation models have shown promise for reducing the data annotation problem and increasing model generalizability and robustness. Especially in the medical image analysis domain where data annotation is considerably more expensive than other fields, and where model robustness is critical, these foundation models hold promise towards increasing performance and adoption of deep learning systems in healthcare. The DINOv2 pre-training approach is specifically promising given its ability to learn general-purpose representations and perform well out-of-the-box in linear probing and kNN settings. We believe that using DINOv2 pre-training on medical data is a promising approach for future research aimed at building large-scale medical foundation models whiteout supervision.
6 Reproducibility
All of the data used for this work is publicly available, and our train, validation, and test split are available here. All the training and validation logs and model weights are available here.
7 Acknowledgments
We thank Prof. Kilian Pohl from Stanford University for the discussion and his suggestions. We also thank King Abdullah International Medical Research Center (KAIMRC) for providing us with NVIDIA RTX A6000 compute nodes.
This study was conducted as part of the Fatima Fellowship program. We thank the Fatima Fellowship and Hugging Face for organizing and sponsoring the Fatima Fellowship Program. [^1] https://www.fatimafellowship.com/
We also thank Meta AI for making the DINOv2 and SAM weights and code bases publicly available.